OpenStudy (anonymous):

Solve for d. Square root d^2 -11=5 1.) 4,-4 2.) 5,-5 3.)5,6 4.)6,-6

OpenStudy (anonymous):

not sure it supposed to be -16 or 16

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

\[\sqrt{d^2-11}=5\] squaring both sides \[d^2-11=25\] \[d^2=25+11=36,d=?\]

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

same.

OpenStudy (anonymous):

It's okay you don't have to answer this I'll try my best to figure it out

OpenStudy (anonymous):

\[d=\pm 6\]

OpenStudy (anonymous):

the answer should be 6,-6

OpenStudy (anonymous):

Okay thanks